Target of Oportunity:
No
Time Observed:
8.0
Time Engineering:
0.0
Time lost technical:
0.25
Time lost weather:
2.0
Time lost other:
0.0
Primary Investigator:
Todd Henry
Email:
Organization:
SMARTS
Total Program Time:
10.25
At 22:00, CHIRON was rebooted because there was communication problems between Nightly observing script website and TCS.
At 23:00 , same problem.It was fixed quickly.
Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
---|---|---|---|---|---|---|
2023-05-14 | 8.0 | 0.0 | 0.25 | 2.0 | 0.0 | 10.25 |
Links
[1] mailto:thenry@astro.gsu.edu
[2] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/42501/delete?destination=print/content/enr-ctio-15-m-2023-05-14
[3] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/42501/edit?destination=print/content/enr-ctio-15-m-2023-05-14
[4] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/add/node/20438?destination=print/content/enr-ctio-15-m-2023-05-14
At the beginning of the night , the dome remained closed because it was cloudy. At 22:05 , the observing program started.
Objects : T0391903064, TOI-3378, CD-547010 and RKS1935-2259 were not observed.